OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 1, -6>; v = <6, 7, -3>; w = <-7, 5, 2>

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

is it cross products?

OpenStudy (jdoe0001):

same way you'd get the dot product for a 2D vector

OpenStudy (anonymous):

so then 29i +9j +79k

OpenStudy (anonymous):

oh...so then it sjust -42+35-6= -13

OpenStudy (jdoe0001):

\(\bf <a, b, c> \cdot <d, e, f> \implies (a\times d)+(b\times e)+(c\times f)\)

OpenStudy (anonymous):

ok, then i got the answer, thanks!

OpenStudy (jdoe0001):

yw